### v1

@@ 2.1 Persistent currents in ultracold atoms
Plotting the ground-state persistent current as a function of the flux threading the system gives a characteristic sawtooth shape with a period fixed by the elementary flux quantum –Figure 2(b). Discontinuous jumps in the current indicate the intersection of two energy parabolas with different angular momenta $\ell$, corresponding to a superposition of these states. This fact should be kept in mind when utilizing the analytical expression, as $\{n\}$ should be changed with increasing flux to get the configuration with the lowest energy, giving the sawtooth current landscape. For fermions depending on the number of particles, the system exhibits different responses to the applied field, causing parabolas with different $\ell$ to traverse each other at $\phi=\frac{j}{2}\phi_{0}$ or $\phi=j\phi_{0}$ for integer $j$ depending on whether $N_{p}$ is odd or even respectively. This so-called parity effect originated as another facet of the Leggett theorem, which utilizes a variational argument to show qualitatively that spinless fermions in a clean ring with arbitrary external potentials and interactions, act as a diamagnet (paramagnet) for odd (even) $N_{p}$ [158] in the presence of the magnetic field - such effects are discussed in further details in Sec. 3.3. Eventually, Loss proved this conjecture in the Luttinger liquid theory framework through the bosonization method [159]. This effect can also be extended to electrons, where it was found that matter-waves of diamagnetic (paramagnetic) character flow in a system consisting of $N_{p}=4\nu+2$ ($N_{p}=4\nu$) particles for $\nu\in\mathbb{Z}^{+}$ and equal population per component [160, 161]. For bosons, there is no parity effect as the current is always diamagnetic. Another noteworthy feature is that the circulation of the particles, be they moving clockwise or anti-clockwise, is captured through the orientation of the persistent current on account of its parity. In the presence of many-body interactions, the persistent current profile, i.e. its magnitude, parity, and periodicity, displays specific dependencies reflecting important aspects of the system. As we will see in Secs. 4 and 5, such a feature makes persistent currents an invaluable diagnostic tool to investigate many-body systems.
@@
The occurrence of persistent currents is not limited to rings in the continuous limit, but can also be extended to lattice systems as well. Performing lattice regularization on the general many-body Hamiltonian (see Eq. (3.1.1) in Sec. 3.1.1) for a system with $N_{s}$ lattice sites, by expanding the field operators in terms of the Wannier functions $\Psi_{\alpha}(\mathbf{r})=\sum_{j}w(\mathbf{r}-\mathbf{r}_{j})a_{j,\alpha}$ [^4] We restricted to the one-band model assuming that the interactions between the bands are weak and to the tight-binding approximation suppressing hoppings beyond nearest-neighbours., we get the tight-binding Hamiltonian $\mathcal{H}_{TB}=-\tilde{J}\sum_{\alpha}\sum^{N_{s}}_{j}(a^{\dagger}_{j,\alpha}a_{j+1,\alpha}+\mathrm{h.c.})$, where the hopping amplitude denoted by $\tilde{J}$ encompasses the flux dependence of the system such that $\tilde{J}=\int w^{*}(\mathbf{r}-\mathbf{r}_{i})\big{[}\frac{1}{2m}(-\imath\hbar\nabla(\mathbf{r})+\chi\mathbf{A}(\mathbf{r}))^{2}+V_{\mathrm{ext}}\big{]}w(\mathbf{r}-\mathbf{r}_{i+1})\mathrm{d}\mathbf{r}$. Going a step further, the gauge field can be incorporated into the Wannier functions allowing that the vector potential is sufficiently smooth on the atomic scale by introducing the transformation $\tilde{w}(\mathbf{r}-\mathbf{r}_{j})=e^{-\imath\Phi(\mathbf{r})}w(\mathbf{r}-\mathbf{r}_{j})\approx e^{-\imath\Phi(\mathbf{r}_{j})}w(\mathbf{r}-\mathbf{r}_{j})$ where $\Phi(\mathbf{r})=\int_{\mathbf{r}_{j}}^{\mathbf{r}}\mathbf{A}(\mathbf{r})\mathrm{d}\mathbf{r}$ with $\mathbf{r}_{j}$ corresponding to the coordinate of lattice site $j$. Consequently, applying this to our model with nearest-neighbour hoppings, the tunnelling parameter acquires a complex phase $\tilde{J}=Je^{\imath\Phi}$ recasting the Hamiltonian into the following form
$\mathcal{H}_{TB}=-\tilde{J}\sum\limits^{N_{s}}_{j=1}(a^{\dagger}_{j}a_{j+1}e^{\frac{2\imath\pi}{N_{s}}\frac{\phi}{\phi_{0}}}+\mathrm{h.c.}).$ (15)
@@
$\mathcal{I}(\phi)=\frac{2\imath\pi J}{\phi_{0}N_{s}}\sum\limits_{j=1}^{N_{s}}(a_{j}^{\dagger}a_{j+1}e^{\frac{2\imath\pi}{N_{s}}\frac{\phi}{\phi_{0}}}-\mathrm{h.c.}),$ (16)
[…]
2.2 Phase coherence, superfluidity and persistent current
@@
Phase coherence is the ability of the quantum wavefunction to retain its phase information. While it is an obvious property to address for single or non-interacting particles, the concept is more subtle for systems of interacting particles [163, 41]. In many-body settings, phase coherence can be quantified in terms of one- and two-body density matrices [10]:
@@
where $\Psi^{\dagger}(x)$ ($\Psi(x)$) are field operators creating (annihilating) a boson or a fermion at position $x$ while we omit vector notations, the arguments in this section apply to any spatial dimensions [^6] All the above relations are readily generalized to lattice systems, by considering the discretized versions of $\rho_{2}$. according to $\rho_{2}(j,l,m,n)=\langle\hat{c}_{j,\uparrow}^{\dagger}\hat{c}_{l,\downarrow}^{\dagger}\hat{c}_{m,\downarrow}\hat{c}_{n,\uparrow}\rangle$. The eigenvectors of $\rho_{1}(x,y)$ define the natural orbitals and the eigenvalues fix their population. Off-Diagonal Long-Range Order (ODLRO) in bosonic systems occurs if one of the eigenvalues is $\lambda_{1}=O(N)$. In this case, the largest eigenvalue of $\rho_{2}(x,x^{\prime},y,y^{\prime})$ is $O(N^{2})$. Instead for fermionic system, ODLRO can occur without a macroscopic occupation of the one-particle density matrix orbital and it s reflected in an $O(N)$ eigenvalue of $\rho_{2}(x,x^{\prime},y,y^{\prime})$. Such a difference reflects the fundamental feature that bosonic systems can indeed Bose-Einstein condense and described by a condensate wave-function $\psi(x)=|\psi(x)|^{2}e^{i\phi(x)}$ that, as we discuss further below, corresponds to the mostly represented natural orbital of $\rho_{1}$; the phase $\phi$ coincides with the single-particle phase. in the case of fermionic systems, the Pauli principle prevents the occupancy of a single quantum level by particles with the same spin, and therefore the many-body coherence needs to be achieved from the single particles’ coherence through higher order correlations [41]. In finite systems or in the presence of interactions, quantum fluctuations can reduce the scaling from $\lambda_{1}(N)$ and $\lambda_{2}(N)$ to $O(N^{\alpha})$ with $10<\alpha<$ 1.
@@
The condensate fraction is defined as $f_{c}=N_{c}/N$, where $N_{c}=|\psi(x)|$ is the effective population of the mostly represented natural orbital of $\rho_{1}$. Notably, the superfluid fraction with $f_{s}$ defines the portion of the system with mass $M_{s}=mNf_{s}$ moving with a superfluid velocity $v(x)=(\hbar/m)\nabla\phi(x)$. We remark that, eventhough the superfluid velocity is expressed through the phase of the condensate $\phi$, $f_{s}$ and $f_{c}$ can be very distinct. A particularly striking example in which $f_{s}$ and $f_{c}$ are indeed very different is provided by 4He at zero temperature in which $f_{s}=N/V$ with a condensate fraction $f_{c}$ that is about the $10\%$ of $N$in a volume $V$ [9, 10].
@@
in which $H_{GI}$ is assumed to be invariant by a local gauge transformation (conserving the total particles number); it may contain interactions and local potentials. We assume that Eq. 19 holds. By relabelling $\psi\rightarrow\psi e^{-i\phi}$, the phase twist can be moved to the canonical momentum:$-i\hbar{{\partial}\over{\partial x}}\rightarrow-i\hbar{{\partial}\over{\partial x}}-\phi$. In this case, it can be proved that
$E_{\phi}-E_{\phi=0}={{1}\over{2m}}+\sum_{\nu\neq 0}{{|\langle\psi_{0}|I|\psi_{\nu}|^{2}}\over{E_{\nu}-E_{0}}},$ (22)
@@
$f_{s}={{2mL^{2}}\over{\phi^{2}\hbar^{2}N}}\left({{1}\over{2m}}+\sum_{\nu\neq 0}{{|\langle\psi_{0}|I|\psi_{\nu}|^{2}}\over{E_{\nu}-E_{0}}}\right).$ (23)
@@
Finally, we note that the superfluid fraction is related to the Drude weight $D_{\omega}\doteq L\pi\left({{\partial^{2}E(\phi)}/{\partial\phi^{2}}}\right)_{\phi=0}$ which is the real part of the linear response to an effective magnetic field [169]. Noting that $\partial^{2}E(\phi)/\partial\phi^{2}$ can be discretized as $(E_{\phi}-E_{\phi=0})/\phi^{2}$, by Eqs. (22) and (23), the Drude weight is related to the superfluid fraction: $D_{\omega}={{\pi\hbar^{2}N}\over{2mL}}f_{s}$. By Eq. (13), $D_{\omega}$ can be calculated as the response of the persistent current to an applied effective magnetic field:
$D_{\omega}=L\pi\left({{\partial I(\phi)}\over{\partial\phi}}\right)_{\phi=0}.$ (24)
In a mesoscopic system, a finite Drude weight reflects a dissipationless flow; normal systems, instead, are characterized by a vanishing Drude weight [170, 171]. Besides being a central property in quantifying the system’s response, the Drude weight has been related to the so-called quantum metric of the ground state [172].
@@ 2.3 Setting cold atoms in motion
Given their charge-neutral nature, cold atoms do not experience a Lorentz force when placed in a magnetic field. However, they are affected electric and magnetic dipole forces in the presence of field gradients, and this makes it possible to engineer an interaction with external fields such that the center of mass motion of an atom mimics that of a charged particle in a magnetic field. While there are still some important practical limits on what can be accomplished with these techniques, there is a large and still-expanding variety of methods for subjecting cold atom systems to artificial gauge potentials [42, 43].
The specific means able to set the condensate in motion heavily depend on the fundamental properties of the order parameter representing the condensate wave function. For the cases in which the order parameter is a simple complex scalar field (as in the bosonic cold atoms), the winding number is conserved as long as the condensate fraction does not vanish. In the cases in which the condensate fraction is characterized by more complicated symmetries (as in fermionic gases), winding numbers can change even without the order parameter presenting any node [158]. The velocity of the matter-wave is not just the gradient of the phase of the condensate wave function but it contains also the gauge field term: $v_{s}=(\hbar/m)(\nabla\phi-\chi A)$ in which $A$ is the artificial vector potential. If the latter is applied sufficiently gradually with $\phi$ staying constant, a flow of matter-wave can be started similarly to what happens for charged particles in magnetic fields. Alternatively, the flow can be start as a specific manipulation of the phase of the condensate wavefunction even without acting on the artificial gauge field. The latter cases correspond to phase imprinting protocols.
Finally, we note that an important difference arises in the nature of the persistent current, depending on the way the effective magnetic field is realized. Specifically, if the motion is started by a stirring laser beam, then the magnetic field is produced as a Coriolis force in the rotating frame of the stirred potential. In this case, the persistent current depends on the stirring rotation displaying a characteristic sawtooth shape in the co-rotating frame-see Sec. 2.1. In contrast, Raman or phase imprinting protocols are able to generate an artificial field in the labframe. Nevertheless, the labframe persistent current versus magnetic field is a staircase in both instances –see Sec. 2.3.

### v3

@@ 2.1 Persistent currents in ultracold atoms
As in the single-particle case, the ground-state persistent current as a function of the flux threading the system gives a characteristic sawtooth shape with a period fixed by the elementary flux quantum –Figure 2(b). Discontinuous jumps in the current indicate the intersection of two energy parabolas with different angular momenta $\ell$, corresponding to a superposition of these states. This fact should be kept in mind when utilizing the analytical expression, as $\{n\}$ should be changed with increasing flux to get the configuration with the lowest energy, giving the sawtooth current landscape. In the presence of many-body interactions, the persistent current profile, i.e. its magnitude, parity, and periodicity, displays specific dependencies reflecting important aspects of the system. As we will see in Secs. 4 and 5, such a feature makes persistent currents an invaluable diagnostic tool to investigate many-body systems.
@@
The occurrence of persistent currents is not limited to rings in the continuous limit, but can also be extended to lattice systems as well. Performing lattice regularization on the general many-body Hamiltonian (see Eq. (3.1.1) in Sec. 3.1.1) for a system with $N_{s}$ lattice sites, by expanding the field operators in terms of the Wannier functions $\Psi_{\alpha}(\mathbf{r})=\sum_{j}w(\mathbf{r}-\mathbf{r}_{j})a_{j,\alpha}$, we get the tight-binding Hamiltonian(1)(1)(1)We restricted to the one-band model assuming that the interactions between the bands are weak and to the tight-binding approximation suppressing hoppings beyond nearest-neighbours. $\mathcal{H}_{TB}=-\tilde{J}\sum_{\alpha}\sum^{N_{s}}_{j}(a^{\dagger}_{j,\alpha}a_{j+1,\alpha}+\mathrm{h.c.})$, where the hopping amplitude denoted by $\tilde{J}$ encompasses the flux dependence of the system such that $\tilde{J}=\int w^{*}(\mathbf{r}-\mathbf{r}_{i})\big{[}\frac{1}{2m}(-\imath\hbar\nabla(\mathbf{r})+\chi\mathbf{A}(\mathbf{r}))^{2}+V_{\mathrm{ext}}\big{]}w(\mathbf{r}-\mathbf{r}_{i+1})\mathrm{d}\mathbf{r}$. Going a step further, the gauge field can be incorporated into the Wannier functions allowing for the vector potential to be sufficiently smooth on the atomic scale by introducing the transformation $\tilde{w}(\mathbf{r}-\mathbf{r}_{j})=e^{-\imath\Phi(\mathbf{r})}w(\mathbf{r}-\mathbf{r}_{j})\approx e^{-\imath\Phi(\mathbf{r}_{j})}w(\mathbf{r}-\mathbf{r}_{j})$ where $\Phi(\mathbf{r})=\int_{\mathbf{r}_{j}}^{\mathbf{r}}\mathbf{A}(\mathbf{r})\mathrm{d}\mathbf{r}$ with $\mathbf{r}_{j}$ corresponding to the coordinate of lattice site $j$. Consequently, applying this to our model with nearest-neighbour hoppings, the tunnelling parameter acquires a complex phase $\tilde{J}=Je^{\imath\Phi}$ recasting the Hamiltonian into the following form
$\mathcal{H}_{TB}=-J\sum\limits^{N_{s}}_{j=1}(a^{\dagger}_{j}a_{j+1}e^{\frac{2\imath\pi}{N_{s}}\frac{\phi}{\phi_{0}}}+\mathrm{h.c.}).$ (11)
@@
$\mathcal{I}(\phi)=\frac{2\imath\pi J}{hN_{s}}\sum\limits_{j=1}^{N_{s}}(a_{j}^{\dagger}a_{j+1}e^{\frac{2\imath\pi}{N_{s}}\frac{\phi}{\phi_{0}}}-\mathrm{h.c.}),$ (12)
which is readily obtained through the Hellmann-Feynmann theorem.(2)(2)(2)$\mathrm{d}E/\mathrm{d}\phi=\langle\Psi(\phi)|\mathrm{d}\mathcal{H}/\mathrm{d}\phi|\Psi(\phi)\rangle$ As we mentioned previously, persistent currents define an instance of the Aharonov-Bohm effect in a closed loop. Such a statement implies that matter-wave currents originate on account of the large coherence length of the circulating particles with respect to the system’s size, i.e. a pure mesoscopic feature [21] (accounted for through the $\frac{1}{N_{s}}$ correction). Furthermore, we point out that the mechanism behind persistent currents in lattice systems, i.e. coherent tunneling, is distinct to that of the continuous regime (hydrodynamics mechanism). Nonetheless, it is known a posteriori that these two mechanisms coincide since in the dilute regime of vanishing lattice spacing, the lattice model tends to its continuous counterpart (see Sec. 3.2.2).
2.2 Phase coherence, superfluidity and persistent currents
@@
Phase coherence is the ability of the quantum wavefunction to retain its phase information. While it is an obvious property to address for single or non-interacting particles, the concept is more subtle for systems of interacting particles [176, 45].
The concept of phase coherence can be discussed phenomenologically through a ‘macroscopic’ wavefunction describing all particles condensed in a single quantum state $\psi_{c}(x)=|\psi(x)|e^{i\theta(x)}$. This picture is relevant either to bosonic condensates or to superconductors at zero temperature. In this case, the expression of the current density in Eq. (7) provides the superfluid velocity $v=j(x,\phi)/|\psi(x)|^{2}$:
$v_{s}=\frac{\hbar}{m}\left(\nabla\theta-\frac{\chi}{\hbar}A\right)$ (13)
and, in turn, a persistent current. We note that the persistent current in systems with a macroscopic phase coherence can also arise because of a suitable phase gradient $\nabla\theta(x)$ of the condensate wavefunction. Such a feature is exploited in ultracold atoms set up through the protocol of phase imprinting (see Sec. 2.3).
In many-body settings, phase coherence can be quantified in terms of one- and two-body density matrices [10]:
@@
where $\Psi^{\dagger}(x)$ ($\Psi(x)$) are field operators creating (annihilating) a boson or a fermion at position $x$ while we omit vector notations, the arguments in this section apply to any spatial dimensions.(3)(3)(3)All the above relations are readily generalized to lattice systems, by considering the discretized versions of $\rho_{2}$ according to $\rho_{2}(j,l,m,n)=\langle\hat{c}_{j,\uparrow}^{\dagger}\hat{c}_{l,\downarrow}^{\dagger}\hat{c}_{m,\downarrow}\hat{c}_{n,\uparrow}\rangle$. The eigenvectors of $\rho_{1}(x,y)$ define the natural orbitals and the eigenvalues fix their population. Off-Diagonal Long-Range Order (ODLRO) in bosonic systems occurs if one of the eigenvalues is $\lambda_{1}=O(N)$. In this case, the largest eigenvalue of $\rho_{2}(x,x^{\prime},y,y^{\prime})$ is $O(N^{2})$. Instead for fermionic system, ODLRO can occur without a macroscopic occupation of the one-particle density matrix orbital and it is reflected in an $O(N)$ eigenvalue of $\rho_{2}(x,x^{\prime},y,y^{\prime})$. Such a difference reflects the fundamental feature that bosonic systems can indeed Bose-Einstein condense and be described by the condensate wave-function $\psi_{c}(x)$ above. As we discuss further below, corresponds to the mostly represented natural orbital of $\rho_{1}$; the phase $\phi$ coincides with the single-particle phase. In the case of fermionic systems, the Pauli principle prevents the occupancy of a single quantum level by particles with the same spin, and therefore the many-body coherence needs to be achieved from the single particles’ coherence through higher order correlations [45]. In finite systems or in the presence of interactions, quantum fluctuations can reduce the scaling from $\lambda_{1}(N)$ and $\lambda_{2}(N)$ to $O(N^{\alpha})$ with $10<\alpha<$ 1.
@@
The condensate fraction is defined as $f_{c}=N_{c}/N$, where $N_{c}$ is the effective population of the mostly represented natural orbital of $\rho_{1}$. Notably, the superfluid fraction with $f_{s}$ defines the portion of the system with mass $M_{s}=mNf_{s}$ moving with a superfluid velocity $v(x)=(\hbar/m)\nabla\phi(x)$. We remark that, even though the superfluid velocity is expressed through the phase of the condensate $\phi$, $f_{s}$ and $f_{c}$ can be very distinct. A particularly striking example in which $f_{s}$ and $f_{c}$ are indeed very different is provided by 4He at zero temperature in which $f_{s}=N/V$ with a condensate fraction $f_{c}$ that is about the $10\%$ of $N$ in a volume $V$ [9, 10].
@@
in which $H_{GI}$ is assumed to be invariant by a local gauge transformation (conserving the total particle number); it may contain interactions and local potentials. We assume that Eq. 16 holds. By relabelling $\psi\rightarrow\psi e^{-i\phi}$, the phase twist can be moved to the canonical momentum:$-i\hbar{{\partial}\over{\partial x}}\rightarrow-i\hbar{{\partial}\over{\partial x}}-\phi$. In this case, it can be proved that
$E_{\phi}-E_{\phi=0}={{1}\over{2m}}+\sum_{\nu\neq 0}{{|\langle\psi_{0}|\mathcal{I}|\psi_{\nu}\rangle|^{2}}\over{E_{\nu}-E_{0}}},$ (19)
@@
$f_{s}={{2mL^{2}}\over{\phi^{2}\hbar^{2}N}}\left({{1}\over{2m}}+\sum_{\nu\neq 0}{{|\langle\psi_{0}|\mathcal{I}|\psi_{\nu}\rangle|^{2}}\over{E_{\nu}-E_{0}}}\right).$ (20)
@@
To close the section, we note that the superfluid fraction is related to the Drude weight that at zero temperature reads $D_{\omega}\doteq L\pi\left({{\partial^{2}E(\phi)}/{\partial\phi^{2}}}\right)_{\phi=0}$ which is the real part of the linear response to an effective magnetic field [167]. Noting that $\partial^{2}E(\phi)/\partial\phi^{2}$ can be discretized as $(E_{\phi}-E_{\phi=0})/\phi^{2}$, by Eqs. (19) and (20), the Drude weight is related to the superfluid fraction: $D_{\omega}={{\pi\hbar^{2}N}\over{2mL}}f_{s}$ – see also [182, 183]. Noting that the first order derivative of the energy is the current $\mathcal{I}$, $D_{\omega}$ can be calculated as the response of the persistent current to an applied effective magnetic field:
$D_{\omega}=L\pi\left({{\partial\mathcal{I}(\phi)}\over{\partial\phi}}\right)_{\phi=0}.$ (21)
In a mesoscopic system, a finite Drude weight reflects a dissipationless flow; normal systems, instead, are characterized by a vanishing Drude weight [184, 185, 186]. Response in systems with spin degrees of freedom (as fermionic ones) can be studied by applying a non-scalar effective magnetic field [187, 184]. Recently, Drude weight in a cigar-shaped cold atom system was measured experimentally [188]. Non-equilibrium transport can be studied by non-linear Drude weights [189, 190, 191]. We finally comment that the Drude weight has been related to the so-called quantum metric of the ground state [192].
@@ 2.3 Setting cold atoms in motion
Given their charge-neutral nature, cold atoms do not experience a Lorentz force when placed in a magnetic field. However, they are affected by electric and magnetic dipole forces in the presence of field gradients, and this makes it possible to engineer an interaction with external fields such that the center-of-mass motion of an atom mimics that of a charged particle in a magnetic field. While there are still some important practical limits on what can be accomplished with these techniques, there is a large and still-expanding variety of methods for subjecting cold atom systems to artificial gauge potentials [46, 47].
The specific means able to set the condensate in motion heavily depend on the fundamental properties of the order parameter representing the condensate wave function. For the cases in which the order parameter is a simple complex scalar field (as in the bosonic cold atoms), the winding number is conserved as long as the condensate fraction does not vanish. In the cases in which the condensate fraction is characterized by more complicated symmetries (as in fermionic gases), winding numbers can change even without the order parameter presenting any node [173]. As we discussed in Sect.2.2, the velocity of the matter-wave is not just the gradient of the phase of the condensate wave function but it contains also the gauge field term: $v_{s}=(\hbar/m)(\nabla\phi-\chi A)$ in which $A$ is the artificial vector potential. If the latter is applied sufficiently gradually with $\phi$ staying constant, a flow of matter-wave can be started as a response to the Aharonov-Bohm flux (see Sec.2.1). Alternatively, the flow can be started as a specific manipulation of the phase of the condensate wavefunction even without acting on the artificial gauge field. The latter cases correspond to phase imprinting protocols.
Finally, we note that an important difference arises in the nature of the persistent current, depending on the way the effective magnetic field is realized. Specifically, if the motion is started by a stirring laser beam, then the magnetic field is produced as a Coriolis force in the rotating frame of the stirred potential. In this case, the persistent current depends on the stirring rotation displaying a characteristic sawtooth shape in the co-rotating frame-see Sec. 2.1. In contrast, Raman or phase imprinting protocols are able to generate an artificial field in the lab frame. Nevertheless, the labframe persistent current versus magnetic field is a staircase in both instances -see Sec. 2.3.